# An Efficient Deep Learning Framework for Brain Stroke Diagnosis Using Computed Tomography (CT) Images

### v2

@@ Abstract
Brain stroke is a leading cause of mortality and long-term disability worldwide, underscoring the need for precise and rapid prediction techniques. Computed Tomography (CT) scan is considered one of the most effective methods for diagnosing brain strokes. Most stroke classification techniques use a single slice-level prediction mechanism, requiring radiologists to manually select the most critical CT slice from the original CT volume. Although clinical evaluations are often used in traditional diagnostic procedures, machine learning (ML) has opened up new avenues for improving stroke diagnosis. To supplement traditional diagnostic techniques, this study investigates machine learning models for early brain stroke prediction using CT scan images. This research proposes a novel machine learning approach to brain stroke detection, focusing on optimizing classification performance with pre-trained deep learning models and advanced optimization strategies. Pre-trained models, including DenseNet201, InceptionV3, MobileNetV2, ResNet50, and Xception, are used for feature extraction. Feature engineering techniques, including BFO, PCA, and LDA, further enhance model performance. These features are then classified using machine learning algorithms, including SVC, RF, XGB, DT, LR, KNN, and GNB. Our experiments demonstrate that the combination of MobileNetV2, LDA, and SVC achieved the highest classification accuracy of 97.93%, significantly outperforming other model-optimizer-classifier combinations. The results underline the effectiveness of integrating lightweight pre-trained models with robust optimization and classification techniques for brain stroke diagnosis.
@@
Machine Learning , Deep Learning , Convolutional Neural Network , Ischemic Brain Stroke , Hemorrhagic Brain Stroke , Hybrid Model , Pre-trained Model
journal: Elsveir
@@ 1 Introduction
Brain Stroke is a medical condition, one of the leading causes of disability and death worldwide, and is brought on by a disruption of blood supply to the brain [1]. Brain stroke occurs when a blockage or ruptured blood vessel interrupts blood flow to the brain, causing oxygen deprivation that damages brain cells. Most of the brain strokes are either ischemic or hemorrhagic [2]. An ischemic stroke occurs when a blood clot or blockage prevents blood flow to the brain, depriving brain tissue of oxygen and nutrients and ultimately resulting in the death of brain cells [3]. A hemorrhagic stroke occurs when a weak blood vessel ruptures, causing an internal or external brain hemorrhage. It is frequently associated with high blood pressure or aneurysms. This kind of stroke directly damages the brain due to the pressure of accumulated blood [4].
@@
Stroke is the second most common cause that contributes significantly to the global mortality burden [5]. Stroke resulted in over 143 million disability-adjusted life years (DALYs) and 6.6 million deaths in 2019 [6]. The age-standardized prevalence rate (ASPR) for stroke was 1,240.263 per 100,000 people. Hemorrhagic strokes had an ASPR of 350.338 per 100,000 people, whereas ischemic strokes, which accounted for around 80% of all stroke occurrences, had an ASPR of 950.973 per 100,000 people [7].
@@
Stroke prevention and prompt response are vital public health priorities because of the tremendous burden this puts on families and healthcare systems throughout the globe. Prevention of brain stroke involves managing risk factors such as high blood pressure, diabetes, obesity, and smoking through lifestyle changes, regular medical check-ups, and medication [8]. Unfortunately, many of the symptoms or risk factors for stroke can be mild or develop gradually, making it difficult to identify those who are at risk, especially in the early stages [9]. The chance of a stroke can be considerably decreased by early identification and suitable therapies, however traditional approaches are reactive rather than predictive and frequently depend on patient-reported symptoms.
@@
Brain stroke can be identified using various conventional methods, which mostly rely on clinical information and patient histories, magnetic resonance imaging (MRI), and computed tomography (CT) scans [10]. In recent years, Machine learning (ML) has emerged as an effective instrument in the diagnosis, treatment, and prognosis of diseases, including stroke. Machine learning can identify brain stroke early using both image data and categorical data [11]. However, many researchers in the past utilized Machine Learning to detect brain stroke by using CT scans, MRI, ECG, or categorical data; for instance, [12] [13] used CT images and deep learning techniques for classifying brain stroke. [14] [15] developed deep learning models to improve early brain stroke identification using MRI images. [16] [17] made use of the Stroke Prediction categorical dataset to detect brain stroke. [18] used raw ECG sensor data to detect stroke at an early stage.
@@
According to the study above, researchers used different kinds of datasets, such as MRI, CT scan, ECG, or Categorical data, to diagnose brain stroke. However, Computed Tomography (CT) Scans and Magnetic Resonance Imaging (MRI) are the most widely used and reliable methods [19]. Deep learning and machine learning hold promise for technological advancement in challenging medical scenarios, especially in brain stroke diagnosis. There are some improvements needed in brain stroke classification using machine learning with CT scan data. This research presents a method that uses a CT scan image dataset to detect brain stroke early and prevent mortality and disability by combining many machine-learning approaches. Our major contributions are as follows.
@@ 2 Literature Review
Numerous researchers have introduced brain stroke detection models using traditional and machine learning techniques. However, this field remains relatively underdeveloped, and ongoing research aims to address its limitations. To gain a clearer understanding of the current state of knowledge in this area, several existing studies that used computed tomography (CT) scans to diagnose brain stroke have been reviewed and presented here for further analysis. A. Abumihsan et al. [20] used a hybrid feature extraction method with a convolutional block attention module (CBAM) to introduce a novel method for ischemic stroke identification from CT scan images. The hybrid feature extraction method provides a thorough representation of brain CT images by leveraging the advantages of two pre-trained models, DenseNet121 and MobileNetV3, through feature fusion. A specialized, first-hand dataset (Dataset 1) collected from the Specialized Private Hospital in Palestine was used to design and assess the proposed methodology. The approach was further tested on a publicly available dataset (Dataset 2) to demonstrate its generalizability and robustness. The findings demonstrate that the suggested model performed exceptionally well, achieving an accuracy of 99.21% on Dataset 1 and 98.73% on Dataset 2. However, this study is limited to only two feature extraction methods, which may restrict feature diversity and model performance. M.M Hossain et al. [21] suggested a novel hybrid ViT-LSTM model to accurately detect stroke using CT scan images. Additionally, they applied Explainable AI (XAI) methods like SHAP, attention mapping, and LIME to provide trustworthy and reasonable predictions and assure clinical relevance. They evaluated the model using BrSCTHD-2023 dataset and the Kaggle brain stroke dataset. The proposed ViT-LSTM model performed the best with 94.55% accuracy with layer normalization and 92.22% accuracy without layer normalization using BrSCTHD-2023 dataset. Besides, it achieved 96.61% accuracy on the Kaggle dataset. Despite this, the hybrid ViT-LSTM model requires higher computational cost, making real-time deployment in clinical settings more challenging. I. Tahyudin et al. [22] developed a deep learning model incorporating Convolutional Block Attention Module (CBAM) with ResNet18 to classify stroke in brain CT images. They used a publicly available dataset containing 2501 brain CT images from patients, with 950 neurological stroke instances and 1551 healthy, non-stroke instances. The CBAM-ResNet18 model outperformed the standard ResNet18 model with 95% validation accuracy. It also obtained an AUC score of 0.99, demonstrating its capability to classify the stroke and non-stroke samples well. Regardless of its outstanding performance, it has some drawbacks, such as generalization issues and limited data size.
@@
SK. UmaMaheswaran et al. [23] presented a novel hybrid technique to detect acute strokes from CT scan images. They used hybrid preprocessing approaches to enhance the quality of images. After that, they used discrete wavelet transform, Gabor Filter, and local binary pattern to extract the features from the segmented images. They then used the dingo optimization algorithm to select optimal features from the extracted images. Afterward, the selected optimal features were input into a novel ML model to detect normal or acute strokes. The proposed XGBoost model achieved an accuracy of 97%. M.A. Saleem et al. [24] developed detection system using CT brain images combining a genetic algorithm to select the most relevant features and a bidirectional long short-term memory (BiLSTM) model to detect brain strokes at the early stage. The BiLSTM model achieved an accuracy of 96.5%. They also used cross-validation to evaluate the model’s overall performance using precision, recall, F1 score, ROC, and AUC. In addition to that, they compared the model’s performance with classical ML models such as SVM, LR, NB, DT, and RF, where the BiLSTM model outperformed all of them. However, the model’s decision-making process is lacking in explainability. M. Sabir et al. [25] introduced a novel deep convolutional neural network (DCNN) model for early-stage brain stroke detection using CT scan images. The model consists of a feature extractor, a feature fusion module, and a stroke detection module. The proposed model performed better than state-of-the-art models like VGG16, ResNet50, and InceptionV3, achieving higher accuracy, sensitivity, and specificity. In their observation, the proposed model has obtained an accuracy of 96.5% to detect stroke within 6 hours of onset. Despite this, the model was trained with limited data, which might have affected its generalizability. S. Prasher et al. [26] developed a deep learning model based on EfficientNet-B0 CNN architecture to predict brain strokes from 2515 CT scan images. The proposed model performed impressively, achieving an accuracy of 98.72%. Thus, they demonstrated its ability to help doctors detect brain strokes from CT images. The researchers think that the model might better identify strokes by being trained with more CT images to provide better diagnoses to stroke patients. However, there is a lack of diversity in the data samples. C.D. Kulathilake et al. [27] presents a multiclass brain stroke detection model using the NCCT dataset, leveraging two deep learning models based on the DenseNet201 transfer learning architecture. The proposed approach follows a two-step process to identify ischemic stroke and its subtypes. To enhance convergence speed and improve training efficiency, the 1-cycle policy and FastAI framework were integrated into the models. Model 01 achieved a training accuracy of 97.03%, a validation accuracy of 94.50%. Similarly, Model 02 attained a training accuracy of 98.02%, a validation accuracy of 95.67%. The DenseNet201 architecture, combined with FastAI techniques, demonstrated superior performance compared to traditional and task-specific algorithms. However, this research experiment used only one architecture and did not explore other architectures. R. Raj et al. [28] proposed a hybrid framework, StrokeViT with AutoML, combining CNNs, Vision Transformers (ViT), and AutoML to improve stroke diagnosis from CT scan images. In this framework, the ViTs were used to extract long-range dependencies, whereas the CNNs were used to capture the local features to improve the slice-wise predictions. AutoML generated patient-level predictions by integrating slice-level data, achieving an accuracy of 87% for the slice-level data as well as successfully classified the patient-level data with 92% accuracy, which surpassed the standalone architectures by an accuracy of 9%. Authors claimed that this technique can adapt for different diseases diagnosed in musculoskeletal and cancer conditions. Nonetheless, the clinical validation is limited in the proposed model. A. Diker et al. [29] introduced a deep learning architecture for categorizing strokes using the Brain Stroke CT scan Dataset. For testing and training purposes, 2501 computed tomography (CT) images of brain strokes were employed in the experimental investigation. This study aimed to categorize brain stroke CT scans as normal and as stroke using a number of well-known pre-trained convolutional neural networks (CNNs), including GoogleNet, AlexNet, VGG-16, VGG-19, and Residual CNN. With an accuracy of 97.06%, VGG-19 obtained the highest score. Though the results are promising, they are limited to a specific and small dataset.
@@
Figure 1: Proposed Architecture for Brain Stroke Detection
@@ 3.1 Dataset Description
The data utilized in this research to detect brain strokes were gathered from a public source and a local clinic. We initially collected two datasets. The name of Dataset 1 is the Brain Stroke CT Image Dataset [30], which contains 2501 images, and the name of Dataset 2 is the brain-stroke-prediction-ct-scan-image-dataset [31], which consists of 2515 images. Together, the two datasets consist of 5016 data. However, Dataset 1 comprises CT scan images into stroke and normal brain data, whereas Dataset 2 consists CT scan images into ischemic and hemorrhagic strokes. Initially, both datasets contained binary classifications. With the help of an expert, we merged the two datasets to construct a three-class dataset comprising Normal, Ischemic, and Hemorrhagic categories. This integration involved incorporating all data from Dataset 2 while selectively include only the normal brain data from Dataset 1. During this process, we removed 424 instances from the original 1,551 normal brain images to enhance the dataset’s quality and relevance. Additionally, we supplemented the dataset by incorporating 177 new normal instances obtained from a local clinic in Bangladesh. The newly generated dataset consists of 3819 CT scan images, split into three categories: Train, Test, and Validation. There are 2397 images in the training set, 919 in the test set, and 503 in the validation set. The three categories of normal, hemorrhagic, and ischemic images are applied to each of these groups showed in Table 2. Samples of each class of brain stroke images are displayed in Fig. LABEL:fig:bi

### v3

# An Efficient Deep Learning Framework for Brain Stroke Diagnosis Using Computed Tomography Images

@@ Abstract
Brain stroke is a leading cause of mortality and long-term disability worldwide, underscoring the need for precise and rapid prediction techniques. A Computed Tomography (CT) scan is considered one of the most effective methods for diagnosing brain strokes. Most stroke classification techniques use a single slice-level prediction mechanism, requiring radiologists to manually select the most critical CT slice from the original CT volume. Although clinical evaluations are often used in traditional diagnostic procedures, machine learning (ML) has opened up new avenues for improving stroke diagnosis. To supplement traditional diagnostic techniques, this study investigates machine learning models for early brain stroke prediction using CT scan images. This research proposes a novel machine learning approach to brain stroke detection, focusing on optimizing classification performance with pre-trained deep learning models and advanced optimization strategies. Pre-trained models, including DenseNet201, InceptionV3, MobileNetV2, ResNet50, and Xception, are used for feature extraction. Feature engineering techniques, including BFO, PCA, and LDA, further enhance model performance. These features are then classified using machine learning algorithms, including SVC, RF, XGB, DT, LR, KNN, and GNB. Our experiments demonstrate that the combination of MobileNetV2, LDA, and SVC achieved the highest classification accuracy of 97.93%, significantly outperforming other model-optimizer-classifier combinations. The results underline the effectiveness of integrating lightweight pre-trained models with robust optimization and classification techniques for brain stroke diagnosis.
@@
Machine Learning , Deep Learning , Convolutional Neural Network , Ischemic Brain Stroke , Hemorrhagic Brain Stroke , Hybrid Model , Pre-trained Model.
journal: arxiv
@@ 1 Introduction
Brain Stroke is a medical condition, one of the leading causes of disability and death worldwide, and is brought on by a disruption of blood supply to the brain [murphy]. Brain stroke occurs when a blockage or ruptured blood vessel interrupts blood flow to the brain, causing oxygen deprivation that damages brain cells. Most of the brain strokes are either ischemic or hemorrhagic [grysiewicz]. An ischemic stroke occurs when a blood clot or blockage prevents blood flow to the brain, depriving brain tissue of oxygen and nutrients and ultimately resulting in the death of brain cells [feske2021ischemic]. A hemorrhagic stroke occurs when a weak blood vessel ruptures, causing an internal or external brain hemorrhage. It is frequently associated with high blood pressure or aneurysms. This kind of stroke directly damages the brain due to the pressure of accumulated blood [montano2021hemorrhagic].
@@
Stroke is the second most common cause that contributes significantly to the global mortality burden [STROKEAHA]. Stroke resulted in over 143 million disability-adjusted life years (DALYs) and 6.6 million deaths in 2019 [feigin2021global]. The age-standardized prevalence rate (ASPR) for stroke was 1,240.263 per 100,000 people. Hemorrhagic strokes had an ASPR of 350.338 per 100,000 people, whereas ischemic strokes, which accounted for around 80% of all stroke occurrences, had an ASPR of 950.973 per 100,000 people [zhang2024global].
@@
Stroke prevention and prompt response are vital public health priorities because of the tremendous burden this puts on families and healthcare systems throughout the globe. Prevention of brain stroke involves managing risk factors such as high blood pressure, diabetes, obesity, and smoking through lifestyle changes, regular medical check-ups, and medication [caplan]. Unfortunately, many of the symptoms or risk factors for stroke can be mild or develop gradually, making it difficult to identify those who are at risk, especially in the early stages [elendu]. The chance of a stroke can be considerably decreased by early identification and suitable therapies, however traditional approaches are reactive rather than predictive and frequently depend on patient-reported symptoms.
@@
Brain stroke can be identified using various conventional methods, which mostly rely on clinical information and patient histories, magnetic resonance imaging (MRI), and computed tomography (CT) scans [ABBASI2023100145]. In recent years, Machine learning (ML) has emerged as an effective instrument in the diagnosis, treatment, and prognosis of diseases, including stroke. Machine learning can identify brain stroke early using both image data and categorical data [daidone]. However, many researchers in the past utilized Machine Learning to detect brain stroke by using CT scans, MRI, ECG, or categorical data; for instance, [patel] [polamuri] used CT images and deep learning techniques for classifying brain stroke. [abulfaraj] [al2022deep] developed deep learning models to improve early brain stroke identification using MRI images. [mondal] [bathla] made use of the Stroke Prediction categorical dataset to detect brain stroke. [choi2021deep] used raw ECG sensor data to detect stroke at an early stage.
@@
According to the study above, researchers used different kinds of datasets, such as MRI, CT scan, ECG, or Categorical data, to diagnose brain stroke. However, Computed Tomography (CT) Scans and Magnetic Resonance Imaging (MRI) are the most widely used and reliable methods [kwong2003computed]. Deep learning and machine learning hold promise for technological advancement in challenging medical scenarios, especially in brain stroke diagnosis. There are some improvements needed in brain stroke classification using machine learning with CT scan data. This research presents a method that uses a CT scan image dataset to detect brain stroke early and prevent mortality and disability by combining many machine-learning approaches. Our major contributions are as follows.
@@ 2 Literature Review
Numerous researchers have introduced brain stroke detection models using traditional and machine learning techniques. However, this field remains relatively underdeveloped, and ongoing research aims to address its limitations. To gain a clearer understanding of the current state of knowledge in this area, several existing studies that used computed tomography (CT) scans to diagnose brain stroke have been reviewed and presented here for further analysis. A. Abumihsan et al. [abumihsan] used a hybrid feature extraction method with a convolutional block attention module (CBAM) to introduce a novel method for ischemic stroke identification from CT scan images. The hybrid feature extraction method provides a thorough representation of brain CT images by leveraging the advantages of two pre-trained models, DenseNet121 and MobileNetV3, through feature fusion. A specialized, first-hand dataset (Dataset 1) collected from the Specialized Private Hospital in Palestine was used to design and assess the proposed methodology. The approach was further tested on a publicly available dataset (Dataset 2) to demonstrate its generalizability and robustness. The findings demonstrate that the suggested model performed exceptionally well, achieving an accuracy of 99.21% on Dataset 1 and 98.73% on Dataset 2. However, this study is limited to only two feature extraction methods, which may restrict feature diversity and model performance. M.M Hossain et al. [hossainmaruf] suggested a novel hybrid ViT-LSTM model to accurately detect stroke using CT scan images. Additionally, they applied Explainable AI (XAI) methods like SHAP, attention mapping, and LIME to provide trustworthy and reasonable predictions and assure clinical relevance. They evaluated the model using BrSCTHD-2023 dataset and the Kaggle brain stroke dataset. The proposed ViT-LSTM model performed the best with 94.55% accuracy with layer normalization and 92.22% accuracy without layer normalization using BrSCTHD-2023 dataset. Besides, it achieved 96.61% accuracy on the Kaggle dataset. Despite this, the hybrid ViT-LSTM model requires higher computational cost, making real-time deployment in clinical settings more challenging.
@@
I. Tahyudin et al. [tahyudin2025high] developed a deep learning model incorporating Convolutional Block Attention Module (CBAM) with ResNet18 to classify stroke in brain CT images. They used a publicly available dataset containing 2501 brain CT images from patients, with 950 neurological stroke instances and 1551 healthy, non-stroke instances. The CBAM-ResNet18 model outperformed the standard ResNet18 model with 95% validation accuracy. It also obtained an AUC score of 0.99, demonstrating its capability to classify the stroke and non-stroke samples well. Regardless of its outstanding performance, it has some drawbacks, such as generalization issues and limited data size. SK. UmaMaheswaran et al. [umamaheswaran2024enhanced] presented a novel hybrid technique to detect acute strokes from CT scan images. They used hybrid preprocessing approaches to enhance the quality of images. After that, they used discrete wavelet transform, Gabor Filter, and local binary pattern to extract the features from the segmented images. They then used the dingo optimization algorithm to select optimal features from the extracted images. Afterward, the selected optimal features were input into a novel ML model to detect normal or acute strokes. The proposed XGBoost model achieved an accuracy of 97%. M.A. Saleem et al. [saleem] developed detection system using CT brain images combining a genetic algorithm to select the most relevant features and a bidirectional long short-term memory (BiLSTM) model to detect brain strokes at the early stage. The BiLSTM model achieved an accuracy of 96.5%. They also used cross-validation to evaluate the model’s overall performance using precision, recall, F1 score, ROC, and AUC. In addition to that, they compared the model’s performance with classical ML models such as SVM, LR, NB, DT, and RF, where the BiLSTM model outperformed all of them. However, the model’s decision-making process is lacking in explainability. M. Sabir et al. [sabir] introduced a novel deep convolutional neural network (DCNN) model for early-stage brain stroke detection using CT scan images. The model consists of a feature extractor, a feature fusion module, and a stroke detection module. The proposed model performed better than state-of-the-art models like VGG16, ResNet50, and InceptionV3, achieving higher accuracy, sensitivity, and specificity. In their observation, the proposed model has obtained an accuracy of 96.5% to detect stroke within 6 hours of onset. Despite this, the model was trained with limited data, which might have affected its generalizability. S. Prasher et al. [prasher2024brain] developed a deep learning model based on EfficientNet-B0 CNN architecture to predict brain strokes from 2515 CT scan images. The proposed model performed impressively, achieving an accuracy of 98.72%. Thus, they demonstrated its ability to help doctors detect brain strokes from CT images. The researchers think that the model might better identify strokes by being trained with more CT images to provide better diagnoses to stroke patients. However, there is a lack of diversity in the data samples. C.D. Kulathilake et al. [kulathilake] presents a multiclass brain stroke detection model using the NCCT dataset, leveraging two deep learning models based on the DenseNet201 transfer learning architecture. The proposed approach follows a two-step process to identify ischemic stroke and its subtypes. To enhance convergence speed and improve training efficiency, the 1-cycle policy and FastAI framework were integrated into the models. Model 01 achieved a training accuracy of 97.03%, a validation accuracy of 94.50%. Similarly, Model 02 attained a training accuracy of 98.02%, a validation accuracy of 95.67%. The DenseNet201 architecture, combined with FastAI techniques, demonstrated superior performance compared to traditional and task-specific algorithms. However, this research experiment used only one architecture and did not explore other architectures. R. Raj et al. [raj2023strokevit] proposed a hybrid framework, StrokeViT with AutoML, combining CNNs, Vision Transformers (ViT), and AutoML to improve stroke diagnosis from CT scan images. In this framework, the ViTs were used to extract long-range dependencies, whereas the CNNs were used to capture the local features to improve the slice-wise predictions. AutoML generated patient-level predictions by integrating slice-level data, achieving an accuracy of 87% for the slice-level data as well as successfully classified the patient-level data with 92% accuracy, which surpassed the standalone architectures by an accuracy of 9%. Authors claimed that this technique can adapt for different diseases diagnosed in musculoskeletal and cancer conditions. Nonetheless, the clinical validation is limited in the proposed model. A. Diker et al. [diker] introduced a deep learning architecture for categorizing strokes using the Brain Stroke CT scan Dataset. For testing and training purposes, 2501 computed tomography (CT) images of brain strokes were employed in the experimental investigation. This study aimed to categorize brain stroke CT scans as normal and as stroke using a number of well-known pre-trained convolutional neural networks (CNNs), including GoogleNet, AlexNet, VGG-16, VGG-19, and Residual CNN. With an accuracy of 97.06%, VGG-19 obtained the highest score. Though the results are promising, they are limited to a specific and small dataset.
Figure 1: Proposed Architecture for Brain Stroke Detection
@@ 3.1 Dataset Description
The data utilized in this research to detect brain strokes were gathered from a public source and a local clinic. We initially collected two datasets. The name of Dataset 1 is the Brain Stroke CT Image Dataset [dataset2501], which contains 2501 images, and the name of Dataset 2 is the brain-stroke-prediction-ct-scan-image-dataset [dataset2515], which consists of 2515 images. Together, the two datasets consist of 5016 data. However, Dataset 1 comprises CT scan images of stroke and normal brain data, whereas Dataset 2 consists of CT scan images of ischemic and hemorrhagic strokes. Initially, both datasets contained binary classifications. With the help of an expert, we merged the two datasets to construct a three-class dataset comprising Normal, Ischemic, and Hemorrhagic categories. This integration involved incorporating all data from Dataset 2 while selectively including only the normal brain data from Dataset 1. During this process, we removed 424 instances from the original 1,551 normal brain images to enhance the dataset’s quality and relevance. Additionally, we supplemented the dataset by incorporating 177 new normal instances obtained from a local clinic in Bangladesh.
@@
(a) Sample images of Computed Tomography (CT) scan in normal, hemorrhagic, and Ischemic states.
The newly generated dataset consists of 3819 CT scan images, split into three categories: Train, Test, and Validation. There are 2397 images in the training set, 919 in the test set, and 503 in the validation set. The three categories of normal, hemorrhagic, and ischemic images are applied to each of these groups, as shown in Table 2. Samples of each class of brain stroke images are displayed in Fig. 19(a)